Karas golf score was one less than three fourths of gregs golf score. If he sum of their scores was 153 find karas golf score
Answer by stanbon(75887)

Karas golf score was one less than three fourths of gregs golf score. If he sum of their scores was 153 find karas golf score
-------
k= (3/4)g - 1
k + g = 153
------------------------
Substitute for "k" and solve for "g"::
(3/4)g - 1 + g = 153
------
(7/4)g = 154
-------------------
(1/4)g = 22
g = 88 (greg's score)
-----
k = (3/4)88 -1 = 65 (Karas's score)
-------------
Cheers,
Stan H.
------------